Wednesday, November 28, 2018 - 08:43

At 6:30 AM we started receiving reports that DHS users are experiencing general network issues.

The appropriate groups have been informed and we are looking to provide a resolution as soon as possible.

This can be tracked through Cherwell Ticket 101024.

[Update]

This has been resolved. Users should be able to use the network as expected.

Sunday, November 25, 2018 - 03:10

Good morning, 

The ADOA-ASET Firewall Migration from the State Data Center at 1510 West Adams Street, in addition to a Layer 3 change performed by AZNet immediately after the Layer 2 Firewall Migration change, were both completed succesfully this morning by 2:42 AM.

This will be the final notification regarding this change.

Thank you, and have a pleasant Sunday.

 

 

Sunday, November 25, 2018 - 00:40

Good morning, 

This morning, starting at 1:00 AM, we will commence with the second attempt at a Next Generation Firewall Migration.

The IBM Operations Team is being contacted to confirm the quiescing of the Mainframe.

We will provide additional updates through the morning until this activity is completed.

Thanks,

Ken Roundtree

 

 

 

Sunday, November 18, 2018 - 05:52

Good morning, 

As reported previously, a routing issue experienced with one of the new Palo Alto firewalls at the Iron Mountain Data Center caused a NO GO decision to be received, and a successful implementation of the Roll Back Plan for this effort was done.

The Firewall Migration will be rescheduled after the root cause of the above mentioned issue has been determined and resolved. 

As of 5:35 AM, the November IPL of the mainframe was completed successfully.

The Technical and Tester Bridge lines have both been terminated at this time.

This will be the final notification for this event.

 

 

 

Sunday, November 18, 2018 - 05:04

Good moring, 

Due to an OSPF routing issue experienced on one of our new Palo Alto firewalls, we received a NO GO decision on moving forward with the Firewall Migration event, and have successfully rolled back to our old firewalls at the State Data Center.   

In the meantime, the IBM Operations team is continuing to move forward with the monthly IPL of the mainframe for November.

The next update will be published at 6:00 AM.

 

Sunday, November 18, 2018 - 04:24

Good morning, 

Unfortaunately, as of 4:20 AM, the routing issue with OSPF on the new secondary firewall could not be resolved.

With this in mind, we have received a NO GO decision, and we are now in the process of implementing our Roll Back Plan.

The Technical Team is still gathering information, and this Firewall Migration effort will be rescheduled for another time in the future.

 

 

 

 

 

Sunday, November 18, 2018 - 03:56

Good moring, 

As of the present time, a routing issue is being worked on by the Technical Team, and the Go/No Go Decision has been pushed back to 4:20 AM. 

Another update will be published at that time.

 

Sunday, November 18, 2018 - 03:03

Good morning, 

At the present time, the firewall migration from the State Data Center to the Iron Mountain Data Center was accomplished successfully.

Now, the applications testers are in the process of testing.

The IBM Operations Team is also in the process of it's monthly IPL for November, and are bringing up the LPARs.

The next update will be published at 4:00 AM.

 

 

Sunday, November 18, 2018 - 01:54

Good morning, 

 

The Mainframe systems were quiesced at 1:30 AM.

At 1:36 AM, ADOA, AZNet and Mosaic technical personnel brought the firewall ports down on the old firewalls at the State Data Center, and brought up the firewall ports on the new firewalls at Iron Mountain Data Center.

Testing, including a failover test, is now in progress.

So far, all looks good.

The next update will be posted at 3:00 AM

 

 

Sunday, November 18, 2018 - 00:51

Good morning -

The outage window for the ADOA-ASET Firewall Migration started at 12:00 AM.

Technical and Tester bridges are currently up and working.

At the present time, the IBM Operations team is performing critical back-ups, prior to bringing down the mainframe for the monthly November IPL.

The current ETA is approximately 1:30 AM before confirmation that all system activity is quiesced, and we will be ready to proceed.

Next Update will be posted at 2:00 AM.

 

 

 

 

Thursday, November 1, 2018 - 15:54

Around 3:45PM we started receiving reports of users within Cisco not being able to answer calls. This is resulting in callers waiting in queue indefinitely. We have alerted the appropriate groups and are looking to provide a resolution as soon as possible.

We are tracking this ticket through Remedy ticket 196087

We apologize for the inconvenience.

Thursday, September 20, 2018 - 13:37

Starting at 1:20PM, started receiving reports of difficulty accessing the YES and azstatejobs.gov and other HR related sites through hr.az.gov. We have notified the appropriate groups and are working to provide a resolution as soon as possible.

This can be tracked through Cherwell ticket 82526.

We apologize for the inconvenience.

Update 2:37PM

Site is back up and reachable without error.

Monday, September 10, 2018 - 12:52

The yes.az.gov website is now working.

The ADOA Service Desk Phone IVR will be updated to reflect this information.

If you experience any issues, you may reach the ADOA Service Desk at 602-364-4444, option 3 to troubleshoot further.

Please reference Cherwell ticket 80108.

Friday, September 7, 2018 - 10:29

Around 10:00AM we started receiving reports of issues connecting to the Copper network. The appropriate group has been notified and is currently working on a solution. There currently isn't an ETA but we look to provide a resolution shortly.

You can track this through Cherwell ticket 79817.

Wednesday, August 8, 2018 - 09:53

The technical issues impacting network connectivity has been resolved

The ADOA Service Desk Phone IVR will be updated to reflect this information.

If you experience any issues, you may reach the ADOA Service Desk at 602-364-4444, option 3 to troubleshoot further.

Please reference Cherwell ticket #72481. 

Wednesday, August 8, 2018 - 09:24

[UPDATE] We are working diligently to repair the outage/network issues and do not currently have an estimated repair time.

We will provide another update at: 10:00AM. 

 

 

Wednesday, August 8, 2018 - 07:33

 

Today, August 8, 2018 we are currently experiencing technical issues with accessing Gmail and internet. We have engaged our Provider to restore services as soon as possible.

This is a known issue, and a resolution is currently being sought.   

At this time, the Estimated Repair time is Unknown.

We apologize for this inconvenience, and hope to be able to report a resolution to this issue shortly.   

Monday, July 30, 2018 - 10:08

At 10:00AM today we started receiving reports of YES's Password Self-Service pages not loading. We have notified the appropriate group and are looking to provide a resolution as soon as possible. We apologize for the inconvenience.

You can follow this issue with Cherwell ticket 70257.

Update:
All traffic to the Self-Service pages is working as expected.

Saturday, July 28, 2018 - 11:36

As of 11:20am all ADOT systems are up and functioning.

If you experience any further issues, please contact (602) 364-4444, Option 2 for our IBM Mainframe support team. 

 

Saturday, July 28, 2018 - 10:56

To All ADOT Online users:

 

ADOT is currently experiencing processing issues with the online systems.

 

ADOT/IBM and ADOA are working to fix the online issues.

 

At this time there is not an estimate of when the corrective action will be completed.

 

Pages